Week of 201026

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Andrew (Nikhef), Borja (Monitoring), Xavier (KIT), Maarten (Alice), Martin Adam (Compute), Andrew (TRIUMF), Dave (FNAL) , Pavlo (ATLAS), Vincent (Security), Xin (BNL), Christian (NDGF), Alberto (Monitoring), Vladimir (LHCb), Julia (WLCG), Pepe (PIC), Darren (NDGF)
  • remote:

Experiments round table:

  • ATLAS reports ( raw view)
    • There’s a known Rucio issue where transfers are submitted with mismatched protocols https://github.com/rucio/rucio/issues/3682. The underlying problem has not yet been identified, but the investigation is ongoing. Such transfers are handled by FTS in streaming mode. This caused further troubles due to the FTS nodes running out of space.

  • CMS reports ( raw view) -
    • CW cannot join the WLCG call today due to a conflicting local meeting
    • Had a number of transfer issues starting early last week involving usually reliable sites (quite some T1s)
      • Investigations point to IPv6 related routing or peering issue in LHCONE and/or LHC-OPN
      • Cause and effects are not fully known by CMS - any insights are welcome
      • Situation much improved towards the end of the working week (actions unclear)
    • Otherwise very good performance: ~220k cores for production + ~55k cores for analysis last week (average)

  • ALICE -
    • NTR

Sites / Services round table:

  • ASGC:
  • BNL: NTR
  • CNAF: NTR
  • EGI: NC
  • FNAL: On 10/21, IPv6 routing problem suspected to be connected to ESNet maintenance earlier in the week, affected connections to CMS global xrootd federation. Workaround by moving to secondary TransAtlantic link. Need to close the loop to see if there was a wider issue as referenced by CW above though.
  • IN2P3: IN2P3-CC will be in maintenance on December 8th, a Tuesday. As usual details will be available one week before the event.
  • JINR:NTR
  • KISTI: NC
  • KIT: ATLAS has reported issues accessing several files that ought to be online all the time (GGUS:149148, GGUS:149151). We could not find out what went wrong so far, but we did identify 27 files, which dCache thought were no longer online, while the replicas actually are found on disk. We've fixed the situation for those files, while we continue investigating the origin of this issue.
  • NDGF: NTR
  • NL-T1: NTR
  • NRC-KI: NC
  • OSG: NC
  • PIC: PIC will be in OUTAGE from Friday 13-11-2020 21:00 [PIC local time] until Sunday 15-11-2020 13:00 [PIC local time]. HTCondor will be stopped at the beginning of the SD.
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services: NTR
  • CERN storage services: NC
  • CERN databases: NTR
  • GGUS: NTR
  • Monitoring:
    • Sitemon ATLAS-CRITICAL flow stopped on Saturday
      • Due to a change in the ETF server providing the "production" sites to process
      • Discussion ongoing with the CRIC team about getting this information directly from there to avoid extra dependencies
    • XrootD migration to MONIT
      • Discussion with CMS/AAA and ATLAS/FAX managers started
      • EOS nodes at CERN restart campaign started this week for ATLAS and CMS
      • We will do an update next week with the remaining nodes.
  • MW Officer: NC
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r24 < r23 < r22 < r21 < r20 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r24 - 2020-10-26 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback